# HaSPeR: An Image Repository for Hand Shadow Puppet Recognition

### v6

@@ Abstract
Hand shadow puppetry, also known as shadowgraphy or ombromanie, is a form of theatrical art and storytelling where hand shadows are projected onto flat surfaces to create illusions of living creatures. The skilled performers create these silhouettes by hand positioning, finger movements, and dexterous gestures to resemble shadows of animals and objects. Due to the lack of practitioners and a seismic shift in people’s entertainment standards, this art form is on the verge of extinction. To facilitate its preservation and proliferate it to a wider audience, we introduce HaSPeR, a novel dataset consisting of 15,000 images of hand shadow puppets across 15 classes extracted from both professional and amateur hand shadow puppeteer clips. We provide a detailed statistical analysis of the dataset and employ a range of pretrained image classification models to establish baselines. Our findings show a substantial performance superiority of skip-connected convolutional models over attention-based transformer architectures. We also find that lightweight models, such as MobileNetV2, suited for mobile applications and embedded devices, perform comparatively well. We surmise that such low-latency architectures can be useful in developing ombromanie teaching tools, and we create a prototype application to explore this surmission. Keeping the best-performing model ResNet34 under the limelight, we conduct comprehensive feature-spatial, explainability, and error analyses to gain insights into its decision-making process. To the best of our knowledge, this is the first documented dataset and research endeavor to preserve this dying art for future generations, with computer vision approaches. Our code and data will be publicly available.
@@ 1 Introduction
Ombromanie, the ancient art of hand shadow puppetry, is a form of art that involves the mesmerizing interplay of light and shadow through the construction and manipulation of shadow figures or silhouettes on a surface, typically a screen or a wall, using one’s hands, body, or props [1, 2]. The alias “cinema in silhouette”111 https://en.wikipedia.org/wiki/Shadowgraphy_(performing_art) is sometimes used to refer to this proto-cinematic medium of entertainment. Its working principle is very straightforward—the puppeteer adeptly positions their hands between a radiant light source and a translucent screen, consequently conjuring shadows and silhouettes that emulate different creatures, as shown in Figure 3.
@@
In tandem with this motivation, this work introduces a seminal addition to the realm of data resources, HaSPeR (Hand Shadow Puppet Image Repository), a methodically curated novel image dataset of hand shadow puppets. The dataset comprises an assemblage of 15,000 samples, that we painstakingly source and verify from 68 professional shadowgraphist clips and 90 amateur shadowgraphist clips. We label and categorize the images with utmost precision to elicit robustness in the image classification models that will undergo training with these images. The samples in HaSPeR are diverse in nature since the source clips are recorded in a plethora of different poses, orientations, and background lighting conditions of the translucent screen. We also inculcate silhouette motion diversity via optical flow estimation [7] in the frame extraction process. We conduct a detailed analysis of HaSPeR’s statistical characteristics. We also employ a variety of state-of-the-art (SOTA) pretrained image classification models to establish a performance benchmark for validating the integrity of the dataset. Additionally, we conduct a thorough evaluation of several facets of the ace ResNet34 model, including its feature representations, interpretability, explainability, and classification errors that it encounters. In an effort to assess the potential of digitized ombromanie teaching tools, we create a simple and lightweight prototype Android application using Flutter for classifying hand shadow puppet images from the phone’s camera feed. We posit that our dataset possesses the potential to offer a wealth of opportunities for exploration and analysis into the artistic domain of hand shadow puppetry.
@@ 4.5 Statistical Analysis
5 Developing Benchmark for HaSPeR
@@ 5.1.1 Baseline Models
For this classification task, we use 31 feature extractor models as baselines, which are listed in Table 2. Some of these models have a track record of good performance across various other image classification tasks [36]. We examine both conventional Convolutional Neural Networks (CNNs) and CNNs augmented with attention mechanisms. Some models have multiple variants in terms of size or number of parameters, and we compare the performance among those variants as well.
@@ 5.1.2 Performance Metrics
We use top-$k$ validation accuracy values (with $k=1,2,3$), Precision, Recall, and F1-score as evaluation metrics to perform comparative analyses among the aforementioned models. The latter three judgment criteria are used due to the slightly imbalanced nature of HaSPeR, as evident in Table 1.
@@ 5.1.4 Hyperparameters and Optimizer
                ShuffleNetV2X10 [38] 2.3M 61.73 78.41 86.1 0.6559 0.6173 0.597 88.73 93.98 96.1 0.8995 0.8873 0.8853
ViTB16 [39] 86.6M 69.71 77.6 83.28 0.7276 0.6972 0.6969 68.88 76.65 81.36 0.7192 0.6868 0.6851
ViTL32 [39] 306.5M 85.1 91.56 94.48 0.872 0.851 0.8509 84.71 91.8 94.08 0.8632 0.8472 0.8465
AlexNet [11] 61.1M 87.01 93.61 95.46 0.884 0.8702 0.8708 88.18 92.58 94.8 0.8887 0.8818 0.8809
SqueezeNet1_1 [40] 1.2M 87.56 92.45 94.15 0.888 0.8757 0.8744 86.21 92.48 94.65 0.8754 0.8622 0.8637
@@
SwinB [42] 87.8M 90.5 95.38 97.4 0.9128 0.905 0.9042 90.2 95.4 97.08 0.9097 0.902 0.9006
GoogLeNet [43] 6.6M 90.73 94.65 95.7 0.9105 0.9073 0.9059 92.18 95.65 96.58 0.9283 0.9218 0.9206
ResNet18 [44] 11.7M 90.91 95.28 96.6 0.9176 0.9092 0.9069 91.25 95.43 97.05 0.9229 0.9125 0.9119
MobileNetV3Large [41] 5.5M 91.2 94.48 95.98 0.9185 0.912 0.911 90.4 94.53 95.26 0.9147 0.904 0.9024
ConvNeXt [45] 88.6M 91.46 96.33 98.05 0.922 0.9147 0.914 92.55 96.36 97.96 0.9306 0.9255 0.9246
SwinV2B [46] 87.9M 91.58 96.25 97.61 0.921 0.9158 0.9151 91.48 96 97.55 0.9209 0.9148 0.9144
VGG16 [12] 138.4M 91.61 95.08 96.65 0.9248 0.9162 0.9168 91 95.21 96.45 0.9235 0.91 0.9119
MNasNet13 [47] 6.3M 91.66 95.65 97.01 0.924 0.9167 0.9149 91.45 95.86 97.26 0.9231 0.9145 0.9133
ConvNextLarge [45] 197.8M 91.88 95.9 97.7 0.9254 0.9188 0.9181 88 94.7 96.56 0.8942 0.88 0.8782
EfficientNetB0 [48] 5.3M 91.93 95.26 96.71 0.9257 0.9193 0.9178 90.4 93.75 95.1 0.9131 0.904 0.9022
MaxViT [49] 30.9M 92.01 96.5 97.81 0.9268 0.9202 0.9214 92.08 95.98 97.36 0.932 0.9208 0.9237
EfficientNetV2S [50] 21.5M 92.31 95.75 96.76 0.9375 0.9232 0.9245 94.45 97.35 98.3 0.9498 0.9445 0.9438
VGG19 [12] 143.7M 92.36 95.13 96.1 0.9354 0.9237 0.9242 91.8 95.06 96.15 0.9296 0.918 0.9187
@@
DenseNet121 [54] 8.0M 92.93 95.75 96.88 0.9367 0.9293 0.9282 92.95 95.51 96.56 0.936 0.9295 0.9285
ResNext101_32X8D [55] 88.8M 93 96.41 97.23 0.9364 0.931 0.9303 94.2 96.61 97.58 0.952 0.942 0.9423
WideResNet101_2 [52] 126.9M 93.36 95.81 96.9 0.9423 0.9337 0.9332 92.73 96.35 97.63 0.9337 0.9273 0.9267
InceptionV3 [56] 27.2M 93.5 96.48 97.35 0.9401 0.935 0.9338 93.71 96.36 97.06 0.9446 0.9372 0.9371
DenseNet201 [54] 20.0M 93.56 95.78 96.73 0.945 0.9357 0.9353 94.43 97 97.61 0.9492 0.9443 0.9442
@@
ResNet34 [44] 21.8M 94.97 97.23 98.23 0.9516 0.9497 0.9491 91.98 95.95 97.2 0.9266 0.9198 0.9189
@@ 5.2.1 Performance Analysis
The ResNet34 model yields the best performance with a top-$1$ accuracy of $94.97\%$. The vanilla version of the model also yields the highest top-$2$ accuracy, top-$3$ accuracy, Precision, Recall, and F1-scores of $97.23\%$, $98.23\%$, $0.9516$, $0.9497$, and $0.9491$ respectively. Upon being equipped with the classifier block shown in Figure 8, the EfficientNetV2S model yields the highest top-$k$ accuracies and Recall. In contrast, the ResNext101_32X8D and DenseNet201 models demonstrate the best performance across Precision and F1-score metrics respectively. At this recess of the performance analysis, we consider the top-$1$ accuracy metric to be the most statistically significant metric. As evident in Table 2, the vanilla models listed in the penultimate row and above lag behind the ResNet34 model when it comes to the top-$1$ accuracy value (as well as the other metrics), which is why we adjudicate that ResNet34 is the best-performing model.
@@ 5.2.4 Practicality Analysis as a Teaching Tool
It is noteworthy to point out that MobileNetV2, with only 3.5 million parameters, managed to surpass many of the other models in terms of performance. This indicates the suitability of this image classification task for lighter, low-latency models that can be used in mobile applications and embedded devices. We create a simple prototype Android application using Flutter to test the efficacy of MobileNetV2 in classifying hand shadow puppet images from the phone’s camera feed. In order to make the prototype work as seamlessly as intended, we make sure the vicinity is well-lit, and the camera accurately captures a sharply focused silhouette. We find that the model has a memory footprint of $29$ MB and achieves an average inference time of $880$ $\mu$s on the Snapdragon 8 Gen 2 mobile chipset featured in the Samsung Galaxy S23 smartphone. Figure 10 portrays the snapshots of the prototype application. There are several other practical implementation challenges involved in this endeavor that we can identify for an educational mobile application to comport well with the target demographics and real-world settings. The system must operate with minimal computational overhead, ensure real-time responsiveness, and maintain low latency. This necessitates a pragmatic tradeoff between the FLOP count and classification accuracy, with the requisite model compression and optimization techniques. Given the potential variability in device camera capabilities, the application must have preprocessing steps including, but not limited to, denoising, adaptive contrast enhancement, and sharpening the input feed to mitigate artifacts. In recognition of the diverse motor capabilities of users, particularly younger learners and users with dexterity impairments, the application must have intelligent motion compensation to stabilize the shaky camera inputs. Adding to the desiderata is the consideration of a suitable UI/UX that is tailored for the pediatric user base, as is done in handwriting teaching apps for children [61], because an intuitive and enjoyable learning experience is of paramount importance for educational apps. This includes providing step-by-step tutorials, interactive guides with progressive difficulty, and illustrative diagrams to demonstrate the creation of hand shadow puppets. A holistically sound approach to this application development endeavor will surely potentiate the pedagogical feasibility of ombromanie.
@@ 6 Conclusion and Future Work
7 Acknowledgments
We convey our heartfelt gratitude, in advance, to the anonymous reviewers for their constructive criticisms and insightful feedback which will surely be conducive to the improvement of the research work outlined in this paper. We also appreciate the Systems and Software Lab (SSL) of the Islamic University of Technology (IUT) for the generous provision of computing resources during the course of this project. Additionally, we wish to acknowledge Shahriar Ivan, Department of Computer Science and Engineering, Islamic University of Technology, for his invaluable assistance in proofreading and offering a preliminary review of this manuscript. We further thank Mohammad Ishrak Abedin and Reaz Hassan Joader, from the same department, for lending their assistance, which was instrumental in refining this paper’s illustrations and diagrams. Syed Rifat Raiyan, in particular, wants to thank his parents, Syed Sirajul Islam and Kazi Shahana Begum, for everything.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data Availability
The codes and datasets generated, written, and/or analyzed during the research project are available in the HaSPeR GitHub repository (https://github.com/Starscream-11813/HaSPeR) and in the HaSPeR Hugging Face repository (https://huggingface.co/datasets/Starscream-11813/HaSPeR).
Funding Sources
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors. All associated costs were self-funded by the authors. Access to computing resources was provided by the Systems and Software Lab (SSL) of the Islamic University of Technology (IUT).

### v7

@@ Abstract
Hand shadow puppetry, also known as shadowgraphy or ombromanie, is a form of theatrical art and storytelling where hand shadows are projected onto flat surfaces to create illusions of living creatures. The skilled performers create these silhouettes by hand positioning, finger movements, and dexterous gestures to resemble shadows of animals and objects. Due to the lack of practitioners and a seismic shift in people’s entertainment standards, this art form is on the verge of extinction. To facilitate its preservation and proliferate it to a wider audience, we introduce HaSPeR, a novel dataset consisting of 15,000 images of hand shadow puppets across 15 classes extracted from both professional and amateur hand shadow puppeteer clips. We provide a detailed statistical analysis of the dataset and employ a range of pretrained image classification models to establish baselines. Our findings show a substantial performance superiority of skip-connected convolutional models over attention-based transformer architectures. We also find that lightweight models, such as MobileNetV2, suited for mobile applications and embedded devices, perform comparatively well. We surmise that such low-latency architectures can be useful in developing ombromanie teaching tools, and we create a prototype application to explore this surmission. Keeping the best-performing model ResNet34 under the limelight, we conduct comprehensive feature-spatial, explainability, and error analyses to gain insights into its decision-making process and explore architectural improvements. To the best of our knowledge, this is the first documented dataset and research endeavor to preserve this dying art for future generations, with computer vision approaches. Our code and data are publicly available at https://github.com/Starscream-11813/HaSPeR.
@@
“Will he not fancy that the shadows which he formerly saw are truer than the objects which are now shown to him?”
Plato, The Republic (Book VII, Allegory of the Cave)
@@ 1 Introduction
Ombromanie, the ancient art of hand shadow puppetry, is a form of art that involves the mesmerizing interplay of light and shadow through the construction and manipulation of shadow figures or silhouettes on a surface, typically a screen or a wall, using one’s hands, body, or props [1, 2]. The alias “cinema in silhouette”333 https://en.wikipedia.org/wiki/Shadowgraphy_(performing_art) is sometimes used to refer to this proto-cinematic medium of entertainment. Its working principle is very straightforward—the puppeteer adeptly positions their hands between a radiant light source and a translucent screen, consequently conjuring shadows and silhouettes that emulate different creatures, as shown in Figure 2.
@@
In tandem with this motivation, this work introduces a seminal addition to the realm of data resources, HaSPeR (Hand Shadow Puppet Image Repository), a methodically curated novel image dataset of hand shadow puppets. The dataset comprises an assemblage of 15,000 samples, that we painstakingly source and verify from 68 professional shadowgraphist clips and 90 amateur shadowgraphist clips. We label and categorize the images with utmost precision to elicit robustness in the image classification models that will undergo training with these images. The samples in HaSPeR are diverse in nature since the source clips are recorded in a plethora of different poses, orientations, and background lighting conditions of the translucent screen. We also inculcate silhouette motion diversity via optical flow estimation [7] in the frame extraction process. We conduct a detailed analysis of HaSPeR’s statistical characteristics. We also employ a variety of state-of-the-art (SOTA) pretrained image classification models to establish a performance benchmark for validating the integrity of the dataset. Additionally, we conduct a thorough evaluation of several facets of the ace ResNet34 model, including its feature representations, feature fusions, interpretability, explainability, and classification errors that it encounters. In an effort to assess the potential of digitized ombromanie teaching tools, we create a simple and lightweight prototype Android application using Flutter for classifying hand shadow puppet images from the phone’s camera feed. We posit that our dataset possesses the potential to offer a wealth of opportunities for exploration and analysis into the artistic domain of hand shadow puppetry.
@@ 4.5 Statistical Analysis
5 Methodology for Benchmarking HaSPeR
@@ 5.1.1 Baseline Models
For this classification task, we use 31 feature extractor models as baselines, which are listed in Table 2. Some of these models have a track record of good performance across various other image classification tasks [36]. We examine both conventional Convolutional Neural Networks (CNNs) and CNNs augmented with attention mechanisms. Some models have multiple variants in terms of size or number of parameters, and we compare the performance among those variants as well. We fuse silhouette-specific features obtained via topological descriptors [37] and polygonization [38] with the extracted features from the best-performing vanilla model (see Appendix 7 for more information).
@@ 5.1.2 Performance Metrics
We use top-$k$ validation accuracy values (with $k=1,2,3$), Precision, Recall, and F1-score as evaluation metrics to perform comparative analyses among the aforementioned models. The latter three judgment criteria are used due to the slightly imbalanced nature of HaSPeR’s professional source clips, as evident in Table 1.
@@ 5.1.4 Hyperparameters and Optimizer
                ShuffleNetV2X10 [40] 2.3M 61.73 78.41 86.10 0.6559 0.6173 0.5970 88.73 93.98 96.10 0.8995 0.8873 0.8853
ViTB16 [41] 86.6M 69.71 77.60 83.28 0.7276 0.6972 0.6969 68.88 76.65 81.36 0.7192 0.6868 0.6851
ViTL32 [41] 306.5M 85.10 91.56 94.48 0.8720 0.8510 0.8509 84.71 91.80 94.08 0.8632 0.8472 0.8465
AlexNet [11] 61.1M 87.01 93.61 95.46 0.8840 0.8702 0.8708 88.18 92.58 94.80 0.8887 0.8818 0.8809
SqueezeNet1_1 [42] 1.2M 87.56 92.45 94.15 0.8880 0.8757 0.8744 86.21 92.48 94.65 0.8754 0.8622 0.8637
@@
SwinB [44] 87.8M 90.50 95.38 97.40 0.9128 0.9050 0.9042 90.20 95.40 97.08 0.9097 0.902 0.9006
GoogLeNet [45] 6.6M 90.73 94.65 95.70 0.9105 0.9073 0.9059 92.18 95.65 96.58 0.9283 0.9218 0.9206
ResNet18 [46] 11.7M 90.91 95.28 96.60 0.9176 0.9092 0.9069 91.25 95.43 97.05 0.9229 0.9125 0.9119
MobileNetV3Large [43] 5.5M 91.20 94.48 95.98 0.9185 0.9120 0.9110 90.40 94.53 95.26 0.9147 0.9040 0.9024
ConvNeXt [47] 88.6M 91.46 96.33 98.05 0.9220 0.9147 0.9140 92.55 96.36 97.96 0.9306 0.9255 0.9246
SwinV2B [48] 87.9M 91.58 96.25 97.61 0.9210 0.9158 0.9151 91.48 96.00 97.55 0.9209 0.9148 0.9144
VGG16 [12] 138.4M 91.61 95.08 96.65 0.9248 0.9162 0.9168 91.00 95.21 96.45 0.9235 0.9100 0.9119
MNasNet13 [49] 6.3M 91.66 95.65 97.01 0.9240 0.9167 0.9149 91.45 95.86 97.26 0.9231 0.9145 0.9133
ConvNextLarge [47] 197.8M 91.88 95.90 97.70 0.9254 0.9188 0.9181 88.00 94.70 96.56 0.8942 0.8800 0.8782
EfficientNetB0 [50] 5.3M 91.93 95.26 96.71 0.9257 0.9193 0.9178 90.40 93.75 95.10 0.9131 0.9040 0.9022
MaxViT [51] 30.9M 92.01 96.50 97.81 0.9268 0.9202 0.9214 92.08 95.98 97.36 0.9320 0.9208 0.9237
EfficientNetV2S [52] 21.5M 92.31 95.75 96.76 0.9375 0.9232 0.9245 94.45 97.35 98.30 0.9498 0.9445 0.9438
VGG19 [12] 143.7M 92.36 95.13 96.10 0.9354 0.9237 0.9242 91.80 95.06 96.15 0.9296 0.9180 0.9187
@@
DenseNet121 [56] 8.0M 92.93 95.75 96.88 0.9367 0.9293 0.9282 92.95 95.51 96.56 0.9360 0.9295 0.9285
ResNext101_32X8D [57] 88.8M 93.00 96.41 97.23 0.9364 0.9310 0.9303 94.20 96.61 97.58 0.9520 0.9420 0.9423
WideResNet101_2 [54] 126.9M 93.36 95.81 96.90 0.9423 0.9337 0.9332 92.73 96.35 97.63 0.9337 0.9273 0.9267
InceptionV3 [58] 27.2M 93.50 96.48 97.35 0.9401 0.9350 0.9338 93.71 96.36 97.06 0.9446 0.9372 0.9371
DenseNet201 [56] 20.0M 93.56 95.78 96.73 0.9450 0.9357 0.9353 94.43 97.00 97.61 0.9492 0.9443 0.9442
@@
ResNet34 [46] 21.8M 94.97 97.23 98.23 0.9516 0.9497 0.9491 91.98 95.95 97.20 0.9266 0.9198 0.9189
                ResNet34 w/ Silhouette Polygonization 21.8M 92.72 96.41 97.51 0.9328 0.9272 0.9257 92.95 $(+1.05\%)$ 95.75 96.61 0.9352 $(+0.93\%)$ 0.9295 $(+1.05\%)$ 0.9283 $(+1.02\%)$
ResNet34 w/ Topological Features 21.8M 93.72 96.43 97.78 0.9432 0.9372 0.9359 94.05 $(+2.25\%)$ 96.45 $(+0.52\%)$ 97.53 $(+0.34\%)$ 0.9476 $(+2.27\%)$ 0.9405 $(+2.25\%)$ 0.9401 $(+2.31\%)$
@@ 5.2.1 Performance Analysis
The ResNet34 model yields the best performance with a top-$1$ accuracy of $94.97\%$. The vanilla version of the model also yields the highest top-$2$ accuracy, top-$3$ accuracy, Precision, Recall, and F1-scores of $97.23\%$, $98.23\%$, $0.9516$, $0.9497$, and $0.9491$ respectively. Upon being equipped with the classifier block shown in Figure 8, the EfficientNetV2S model yields the highest top-$k$ accuracies and Recall. In contrast, the ResNext101_32X8D and DenseNet201 models demonstrate the best performance across Precision and F1-score metrics respectively. At this recess of the performance analysis, we consider the top-$1$ accuracy metric to be the most statistically significant metric. As evident in Table 2, the vanilla models listed in the upper part’s penultimate row and above lag behind the ResNet34 model when it comes to the top-$1$ accuracy value (as well as the other metrics), which is why we adjudicate that ResNet34 is the best-performing model. We hypothesize that residual connections in ResNets help preserve low-level edge and contour information through identity mappings, ensuring that crucial silhouette boundaries aren’t lost as the network deepens, making them better at capturing subtle variations.
@@ 5.2.4 Practicality Analysis as a Teaching Tool
It is noteworthy to point out that MobileNetV2, with only 3.5 million parameters, managed to surpass many of the other models in terms of performance. This indicates the suitability of this image classification task for lighter, low-latency models that can be used in mobile applications and embedded devices. We create a simple prototype Android application using Flutter to test the efficacy of MobileNetV2 in classifying hand shadow puppet images from the phone’s camera feed. In order to make the prototype work as seamlessly as intended, we make sure the vicinity is well-lit, and the camera accurately captures a sharply focused silhouette. We find that the model has a memory footprint of $29$ MB and achieves an average inference time of $880$ $\mu$s on the Snapdragon 8 Gen 2 mobile chipset featured in the Samsung Galaxy S23 smartphone. Figure 10 portrays the snapshots of the prototype application. There are several other practical implementation challenges involved in this endeavor that we can identify for an educational mobile application to comport well with the target demographics [63] and real-world settings. The system must operate with minimal computational overhead, ensure real-time responsiveness, and maintain low latency. This necessitates a pragmatic tradeoff between the FLOP count and classification accuracy, with the requisite model compression and optimization techniques. Given the potential variability in device camera capabilities, the application must have preprocessing steps including, but not limited to, denoising, adaptive contrast enhancement, and sharpening the input feed to mitigate artifacts. In recognition of the diverse motor capabilities of users, particularly younger learners and users with dexterity impairments, the application must have intelligent motion compensation to stabilize the shaky camera inputs. Adding to the desiderata is the consideration of a suitable UI/UX that is tailored for the pediatric user base, as is done in handwriting teaching apps for children [63], because an intuitive and enjoyable learning experience is of paramount importance for educational apps. This includes providing step-by-step tutorials, interactive guides with progressive difficulty, and illustrative diagrams to demonstrate the creation of hand shadow puppets. A holistically sound approach to this application development endeavor will surely potentiate the pedagogical feasibility of ombromanie.